# NICE: To Optimize In-Context Examples or Not?

### v1

@@ 3.2 Normalized Invariability to Choice of Examples (NICE)
where $S(j,\text{I})$ is the score for the $j$-th bin in $\mathcal{B}$ for a given instruction. Note that this score can be computed for any choice of task, any instruction, any performance measure $Y_{LM}$, and any grouping function $f$ for partitioning the candidate examples. Liu et al. (2021) show that cosine similarity with the query is a good proxy for selecting in-context examples. Thus, we set up our baseline to study the performance trend in various datasets by binning the examples ($\mathbf{x}$) based on their similarity score with the query ($\mathbf{q}$). We assume that the grouping function $f$ to be as follows:
@@
We partition the candidate examples into bins based on the value of the grouping function $f$ for example $\mathbf{x}$, e.g., $0-10\%$, $10-20\%$. Combining Eqs. (1) and (2) with the grouping function from Eq. (3) gives us the NICE baseline we set up to distinguish tasks.
We hypothesize that for tasks with a high NICE, the task is learnable from the instructions and does not require carefully selected examples. On the other hand, for tasks with a lower NICE, examples play a major role in task learning, so we suggest spending more budget on optimizing ICE. A simpler task (such as SST-2) can be expected to have a flat trend in performance across bins, and thus a high NICE. On the other hand, for complex tasks where the model needs to understand the input-output relationship and its syntax (such as MTOP, Break and NL2BASH in our study) by looking at the retrieved in-context examples, we expect to have a lower NICE. Our simple baseline using cosine similarity-based bins can be easily extended to consider more advanced clustering strategies by using heuristics such as influence functions (Nguyen and Wong, 2023).
We experiment with the following datasets in line with other works in in-context learning: Stanford Sentiment Treebank (SST-2 and SST-5) (Socher et al., 2013), question multi-class classification in TREC (Li and Roth, 2002), textual entailment in MNLI (Williams et al., 2017), semantic parsing in MTOP which is a multilingual task-oriented semantic parsing dataset with hierarchical intent-slot schema.(Li et al., 2020), question decomposition in Break where input is a complex question (x) and the task is to map the question to its QDMR steps ($\mathbf{s}=\langle s_{1},...,s_{n}\rangle$) (Wolfson et al., 2020), code generation in NL2BASH, and hybrid and multi-choice question-answering in FinQA (Chen et al., 2021) and MedQA (Jin et al., 2021). We download standard versions of all these datasets from HuggingFace (Delangue and team., 2016). Details about our tasks along with their NICE with and without instructions are given in Table 1.
Figure 2: Per-bin performance trend of various datasets with distance-based bin using GPT-4. Arrow indicates change in NICE when shifting from no instruction to a detailed instruction. Note that the y-axis shows the ratio of performance of a given bin with the max performance over all bins with no instructions.
@@
Figure 3: Comparison of in-context example optimization methods against instructions for various classification and structured generation tasks using GPT-3.5-turbo. Abbreviation key: LM (Task Def. with Label Meanings), R (Task Def. with Rules). All ICL methods pick $8$-shot examples.
4 Experiment Design
4.1 Prompt Structure and Instruction Set
For a fair comparison across instructions and easy applicability to any new task, we use a standard prompt template that consists of an instruction, $4$-shot demonstrations $(x_{i},y_{i})$ from a training set $\mathcal{D_{T}}$, and the query $x$ from the test set with a cue for the model to predict the corresponding $y$. We define a consistent structure to create a general set of instructions for all tasks, iteratively adding useful, task-specific information to the instruction starting with no instruction following a guideline similar as in Zhang et al. (2023). A consistent template that does not involve human creativity ensures that our findings are not constrained to a narrow distribution of carefully-crafted prompts. We use this gradation of adding detailed information as a rough proxy for "improving" the instruction based on results from Zhang et al. (2023).
  • No Instruction (NI): We provide no instruction and the model is expected to learn only from in-context examples.
  • Task Definition with Label Space (+LS): Including the label space in the task definition. (e.g. [“positive”, “negative”] for a binary sentiment classification task.
  • Delusive Instruction (DI): We also evaluate the model’s task learnability when the instruction is misleading, i.e., describing a task with shuffled labels. (e.g. “positive” for “negative”).
For generative tasks, we experiment with the ‘No-Instruction’ and ‘Task-Definition’ cases in the same way as defined for classification tasks. Since there is no notion of label space in generation tasks, we introduce an additional variety of detailed instruction:
  • Task Definition with Rules: A simple task definition accompanied with syntactic rules for the task.
@@ 4.1 Prompt Structure and Instruction Set
Figure 4: Label choice ablation (gold to random) performance for classification and structured generation tasks using GPT-3.5-turbo. Abbreviation key = DI (Delusive Instruction), NI (No Instruction), TD (Task Definition), +LS (Task Def. with Label Space), +LM (Task Def. with Label Meanings), +R (Task Def. with Rules). Note that classification tasks do not have rules while generation tasks do not have label spaces and delusive instructions.
4.2 Investigations
We structure our study around the following research questions:
RQ1: What kinds of tasks require in-context example optimization?
To investigate whether ICE optimization is needed for a given task, we evaluate various tasks based on the NICE metric defined in Eq. (2).
@@ 4.2 Investigations
Setup: We assume that the training split of each dataset acts as a pool of candidate examples, and pick 50 queries randomly from the test split and partition the training set for each query into groups/bins based on the grouping function $f$ defined in Eq. (1). We randomly sample 10 sets of 4-shot demonstrations from a specific bin for a given query and evaluate the model’s performance and report the averaged out results over 50 queries for each group. For better visualization, we scale the accuracies across the different groups by dividing them by their mean. This scaling of accuracy values doesn’t affect the trend as our metric is invariant to linear scaling (see Subsection 3.1).
RQ2: How do different kinds of in-context example selection approaches fare against a detailed instruction?
@@
We consider three different baselines for methods to select in-context examples for a give query based on different kinds of properties:
@@
Top-$k$-BERT: This dense retrieval method uses BERT (Devlin et al., 2018) embeddings to compute the cosine similarity of examples with the query and selects the top-$k$ examples.
@@
DPP: Determinantal Point Processes (DPP) Borodin (2009) computes a probability measure over the possible subsets of a discrete set. They model the repulsion between the items in a chosen subset by considering the volume of the parallelepiped formed by the feature vectors of the items in the chosen subset as a measure of probability of picking up that subset, as of consequence of which, the likelihood of selecting a subset decreases as items become more similar. Therefore, DPP ensures that the chosen subset is diverse in terms of specified criteria, such as representativeness.
Using each of the above methods, we pick up the top $k=8$ examples as the in-context examples with the instruction, similar to other works that use $k\in\{4,8,16\}$ (Min et al., 2022a; Wei et al., 2023; Lu et al., 2021; Zhang et al., 2022b).
RQ3: How much do the ground-truth labels of ICE matter as we keep on improving the instruction based on a fixed template?
Min et al. (2022b) identify four aspects of demonstrations which serve as a learning signal to the LLM: the input-label mapping, the distribution of the input text, the label space and the format. Concretely, they say that in-context examples play a role by informing the model of the label space or the format of the output. Pan et al. (2023) discuss the phenomena of task learning and task recognition as the two ways through which an LLM learns from demonstrations by comparing the performance of randomly labeled examples against abstract and gold-labeled ones.
Relatedly, in this work, we evaluate the role of input-label mapping for in-context learning. Specifically, we investigate whether, instead of relying on the demonstrations for information about the label space or format of the output, it is possible to subsume this information in the instruction such that this leads to a minimal drop in performance. To this extent, we introduce perturbations in the in-context examples by randomizing the ground-truth labels of the in-context examples. In the case of classification tasks, we assign randomly chosen labels $y\sim\mathcal{Y_{T}}$ to the inputs of the in-context examples, where $\mathcal{Y_{T}}$ denotes the label space of the output for a particular task $\mathcal{T}$. In the case of generation tasks, we shuffle the ground truth values $y$ of the in-context examples amongst themselves.
5 Results
We devise three experiments to answer the three research questions highlighted in 4.2 and present the results for them in this section. We experiment with state-of-the-art LLMs with a good instruction following ability, and use GPT-4-Turbo for computing the NICE score for each task, and GPT-3.5-Turbo for all other experiments. We also show NICE results from open-source models in Appendix A.3.
5.1 High-NICE and Low-NICE Tasks (RQ1)
  1. 1.
    No Instruction: When we prompt the model with just in-context examples without any instruction, we observe the NICE for relatively simpler tasks (SST2, SST5, and MNLI) to be higher than more complex tasks (TREC, Break, NL2BASH, and MTOP).
  2. 2.
    Detailed Instruction: When a detailed task-specific instruction is added to the prompt along with in-context examples, we observe a flattening effect of the instructions for TREC, SST5, and MTOP, thus increasing the NICE for these tasks. This evidence supports our hypothesis that the effect of in-context example optimization decreases in the presence of instructions.
We observe in Fig. 2 that for tasks with a high NICE, it is not helpful to selectively retrieve in-context examples conditioned on the query.
As described in 2, Rubin et al. (2021) propose a contrastive-learning objective for selecting in-context examples. Specifically, they define a score function for each of the examples in the set of candidates given by:
$s_{ij}=P_{LM}(y_{i}|e_{ij},x_{i})\quad\forall j\in\{1,2,...,L\}\quad\forall i$
In their case, the contrastive learning objective plays the role to increase the relevance of the subset of examples with higher LM scores over the others for a given query. We show that in a scenario, where we have example equivalence in terms of model’s performance across the different candidate examples, even this exercise doesn’t lead to an improvement in performance.
5.2 Optimizing In-context Examples: A Comparison with Instructions (RQ2)
@@
On the other hand, for tasks with a low NICE (MTOP, Break, and NL2BASH), we see that in-context example optimization helps in improving the performance over randomly sampled examples accompanied with instructions. Interestingly, we observe that the the distance of the task’s NICE with complete instruction learnability (i.e., NICE =1) is proportional to the difference in performance of in-context example optimization and instructions. For instance, for tasks such as NL2BASH and SST5, we see a slight dip in performance when using instruction, while the dip is significant for MTOP and Break.
5.3 Label Perturbation (RQ3)
@@
In Fig. 4, we see that as we move to more detailed instructions, the difference in performance with gold-labeled examples against randomly-labeled examples keeps on diminishing and ultimately becomes negligible for detailed task-specific instructions. This validates our hypothesis that for tasks with a high NICE, good instructions suffice and the dependence of the performance on examples is less and it reduces further with better instructions. (see 4.1 for the instruction templates). Thus, we see that a better instruction helps in reducing the model’s reliance on the in-context examples to understand input-label mapping. In the case of classification tasks, there is a drop in performance with the incorporation of a delusive instruction in the prompt; this drop is also an indicator of the model’s dependence on the instruction to understand a task.
@@
Finally, on tasks with a low NICE (MTOP, Break and NL2BASH), we see a different trend in Fig. 4. While with no instructions we see a significant improvement in the performance ($\sim$1.04x for MTOP and $\sim$1.25x for Break), the improvement from gold-labeled examples stays consistent with the addition of task definitions and the syntax rules of the task. Furthermore, the narrowing of the performance gap between gold and random labels for in-context examples can also be observed in the case of these tasks, albeit at a slower rate compared to tasks with a high NICE. Consequently, the importance of optimizing in-context examples becomes more evident in this scenario.
@@ 6 Conclusion
In this work, we challenge the assumption that optimizing in-context examples is always crucial for better performance of LLMs on downstream tasks. Through extensive experimentation on instruction-tuned models, we demonstrate empirically that ICE optimization yields diminishing returns in the presence of task-specific instructions for a class of tasks. We introduce Normalized Invariability to Choice of Examples (NICE) as a metric to analyze the learnability of tasks from given instructions, and show that for hig-NICE tasks, better-quality instructions lead to comparable or better performance than ICE optimization baselines. We also conduct label choice ablation experiments, confirming that with detailed instructions, the model’s performance becomes less sensitive to the choice of labels in the in-context examples. We hope that our work sheds some light on the effects of instructions and examples for prompting LLMs and would lead to a more systematic method to optimize prompts for in-context learning of any new task.
@@ A.1.1 Classification Tasks
In 4.1, we discuss the structured instruction sets we use. In this section, we give some concrete examples. We also note that all of these instructions follow standard templates and none of them required human creativity. The total amount of time it took to create all of them for $9$ datasets was under $1$ hour. We show the instruction sets for SST-2 and MTOP, and refer to our codebase for all other datasets.
@@ A.1.2 Generation Tasks
3. The standard format of the output is:- [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]…[SL: <slot-label-n> <slot-label-value-n> ] ] if there are n slot labels tagged in an utterance.
4. In some cases, the slots can be nested with intents within them, for those cases, use this format:- [IN: <user-intent> [SL: <slot-label> [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]…[SL: <slot-label-n> <slot-label-value-n>]]]]
@@
The “No Instruction” case for both classification and generation tasks has an empty instruction appended by the examples. For instruction templates for all our classification and generation tasks, we refer the interested reader to our codebase.

### v3

@@ 4.2 Normalized Invariability to Choice of Examples (NICE)
where $S(j,\text{I})$ is the score for the $j$-th bin in $\mathcal{B}$ for a given instruction. Note that this score can be computed for any choice of task, any instruction, any performance measure $Y_{LM}$, and any grouping function $f$ for partitioning the candidate examples. Liu et al. (2022) show that cosine similarity with the query is a good proxy for selecting ICE. Thus, we set up our baseline to study the performance trend in various datasets by binning the examples ($\mathbf{x}$) based on their similarity to the query ($\mathbf{q}$). We assume the grouping function $f$ as,
@@
We partition the set of examples into bins based on the grouping function $f$ for each example $\mathbf{x}$, e.g., $0-10\%$, $10-20\%$. Combining Eqs. (1) and (2) with the grouping function from Eq. (3) gives us the NICE baseline we set up to distinguish tasks. Note that our simple cosine-similarity-based baseline can be easily extended to consider advanced clustering heuristics such as influence functions (Nguyen and Wong, 2023).
We hypothesize that high-NICE tasks are learnable from instructions and do not need carefully selected examples, whereas for low-NICE tasks, examples play a major role and we suggest spending more budget on optimizing ICE. We validate this hypothesis in the next section.
Figure 2: GPT-4’s performance trend on various tasks as the average distance of ICE from the query increases. Arrow indicates change in NICE from no instruction to a detailed instruction. With a detailed instruction, NICE is near 1 for all tasks except schema-based generative tasks like MTOP and Break. Error bars are calculated from the standard deviation of accuracies across 50 queries.
@@
Figure 3: Comparing ICE optimization methods against random ICE with a detailed instruction for classification and structured generation tasks using GPT-3.5. LM (Task Def. with Label Meanings), R (Task Def. with Rules).
@@
Figure 4: Label perturbation (gold to random) performance for classification and structured generation tasks using GPT-3.5-turbo. Abbreviation key = DI (Delusive Instruction), NI (No Instruction), TD (Task Definition), +LS (Task Def. with Label Space), +LM (Task Def. with Label Meanings), +R (Task Def. with Rules). Error bars are calculated from the standard deviation in accuracies in the same way as in Fig. 3 (random example case).
5 Results
5.1 Experimental Setup
To calculate the NICE metric, we use GPT-4-Turbo and due to cost considerations, we use GPT-3.5-Turbo for all other experiments. We show similar results using GPT-3.5-Turbo and the open-source models Llama-2-70B-Chat and Mixtral-8x7B-Instruct in Appendix A.4.
Using the ICE baseline methods from Subsection 3.2, we pick up the top $k=8$ examples as the in-context examples with the instruction, similar to other works that use $k\in\{4,8,16\}$ (Min et al., 2022; Wei et al., 2023; Lu et al., 2022; Zhang et al., 2022c).
Tasks: We experiment with the following tasks in line with other works in in-context learning: Stanford Sentiment Treebank (SST-2 and SST-5) (Socher et al., 2013), question multi-class classification in TREC (Li and Roth, 2002), textual entailment in MNLI (Williams et al., 2017), semantic parsing in MTOP, which is a multilingual task-oriented semantic parsing dataset with hierarchical intent-slot schema (Li et al., 2021), question decomposition in Break, where the input is a complex question (x) and the task is to map the question to its QDMR steps ($\mathbf{s}=\langle s_{1},...,s_{n}\rangle$) (Wolfson et al., 2020), code generation in NL2BASH (Lin et al., 2018) and hybrid and multi-choice question-answering in FinQA (Chen et al., 2021) and MedQA (Jin et al., 2021) respectively. Details about our tasks and their NICE scores are given in Table 1.
Train-test split: The training split of each dataset acts as a pool of candidate examples, and pick 50 queries randomly from the test split as our tet set.
@@ 5.1 Experimental Setup
Computing NICE: For computing the NICE metric, for each test query, we partition the candidate examples into groups/bins based on the grouping function $f$ defined in Eq. (3). We randomly sample 10 sets of k-shot demonstrations  [^1] For GPT-4-Turbo and GPT-3.5-Turbo models, we take k=4 while we set k=16 for the others. from a specific bin for a given query and evaluate the model’s performance and report the averaged out results over 50 queries for each group.
5.2 RQ1: Does ICE optimization always help?
  • No Instruction: When we prompt the model with just in-context examples without any instruction, the NICE for relatively simpler tasks (SST2, SST5, and MNLI) is higher than more complex tasks (TREC, Break, NL2BASH, and MTOP).
  • Detailed Instruction: When a detailed, task-specific instruction is added to the prompt along with ICE, we observe a flattening effect owing to improved invariability to bin-selection in bin-wise performance for TREC, SST5, and MTOP, thus increasing the NICE for these tasks. This evidence supports our hypothesis that the effect of picking specific ICE decreases in the presence of instructions.
We observe in Fig. 2 that for tasks with a high NICE, it is not helpful to selectively retrieve in-context examples conditioned on the query. Learning-based methods such as Rubin et al. (2022) propose a contrastive-learning based objective for selecting ICE by scoring each subset of examples for a given query. With invariability in terms of the model’s performance across appropriately grouped subsets, such scoring-based methods would also not lead to performance improvements.
5.3 RQ2: Optimized ICE versus Random ICE
@@
We consider three different baselines for methods to select in-context examples for a given query.
@@
  • Top-$k$-BERT: This dense retrieval method uses BERT (Devlin et al., 2019) embeddings to compute the cosine similarity of examples with the query and selects the top-$k$ examples.
@@
  • DPP: Determinantal Point Processes (DPP) Borodin (2009) compute a probability measure over the possible subsets of a discrete set. They model the repulsion between the items in a chosen subset by considering the volume of the parallelepiped formed by the feature vectors of the items in it as a measure of the probability of picking up that subset, thus optimizing for diversity in a specified criterion.
To ensure the reliability of our results, we report the mean accuracies with standard deviation bars. To obtain the error bar for the Instruction + Random Examples approach, we sample 3 random subsets of $k$ in-context examples. We then compute the mean and standard deviation of accuracy for each query and aggregate these mean accuracies and standard deviations across different queries in the dev set of each dataset. We report the mean of these mean accuracies (for different dev-set queries) and the mean of these standard deviations (serving as the error bar). For the ICE selection methods (DPP, Top-K-BERT, BM25), we consider 3 different permutations of the in-context examples retrieved by these methods for each query. We report the final mean and standard deviation in accuracy similar to the random example case above.
@@
On the other hand, for tasks with a low NICE (MTOP, Break, and NL2BASH), we see that ICE optimization helps in improving the performance over randomly sampled examples accompanied with instructions. Interestingly, we observe that the the distance of the task’s NICE with complete instruction learnability (i.e., NICE =1) is proportional to the difference in performance of in-context example optimization and instructions. For instance, for NL2BASH and SST5, we see a slight dip in performance when using instruction, while the dip is significant for MTOP and Break.
5.4 RQ3: Do ground-truth labels matter?
@@
In Fig. 4, we see that as we move to more detailed instructions, the difference in performance with gold-labeled examples against randomly-labeled examples keeps on diminishing and ultimately becomes negligible for detailed task-specific instructions. This validates our hypothesis that for tasks with a high NICE, good instructions suffice and the dependence of the performance on examples is less and it reduces further with better instructions. (see 3.1 for the instruction templates). Thus, we see that a better instruction helps in reducing the model’s reliance on the in-context examples to understand input-label mapping. In the case of classification tasks, there is a drop in performance with the incorporation of a delusive instruction in the prompt; this drop is also an indicator of the model’s dependence on the instruction to understand a task.
@@
Finally, on tasks with a low NICE (MTOP, Break and NL2BASH), we see a different trend in Fig. 4. While with no instructions we see a significant improvement in the performance ($\sim$1.04x for MTOP and $\sim$1.25x for Break), the improvement from gold-labeled examples stays consistent with the addition of task definitions and the syntax rules of the task. Furthermore, the narrowing of the performance gap between gold and random labels for in-context examples can also be observed in these tasks, albeit at a slower rate compared to tasks with a high NICE. Consequently, the importance of optimizing in-context examples becomes more evident.
6 Discussion
6.1 Which tasks are sensitive to ICE choice?
Our results indicate that tasks that require the output to follow complex schema, such as MTOP, Break and NL2BASH, have a low NICE score. As a result, they are more sensitive to ICE optimization than other tasks. For example, in the MTOP dataset (shown in Fig. 11), the output format can vary depending on the query. Thus, LLMs may require carefully selected in-context examples tailored to specific queries to output the correct format.
To further explore the role of output schema in the sensitivity of a task to ICE, we present two ablation experiments where we vary the output schema complexity of tasks. For the tasks with a standard output format such as SST2 and SST5, we add a transformation that induces a novel schema/structure in the output. We find that the NICE score decreases significantly when the output has a non-standard, complex structure. More details can be found in A.6.1. As another experiment, we use the GSM-8k Cobbe et al. (2021) dataset. We consider two kinds of required output schema: 1) Final answer correctness: We only consider the last numerical value in the LLM’s response and compare against the ground-truth; 2) Reasoning format correctness: We use regular expressions to check if the model adheres to the reasoning syntax. Consistent with our hypothesis, we find that the second task (with complex output format) has a lower NICE score than the first task. Please refer to A.6.2 for the details and results of this experiment.
6.2 Key Takeaways
Through extensive experimentation, we found that for most tasks, except those with complex requirements on output format (which can be query-dependent, see 6.1), optimizing in-context examples did not significantly impact performance once detailed instructions were provided. For simpler tasks, random ICE with detailed instructions performed as well as optimized ICE.
Tasks with complex output formats, such as Break, MTOP, NL2BASH (see Tab. 1, Fig. 2), SST2/SST5 (with induced schema) and GSM-8k (format-correctness) require ICE optimization (see Appendix  A.6.1 and  A.6.2). Our metric indicates the necessity of ICE optimization in these cases, as the NICE score is low. For these tasks, dynamically selecting ICE helps the LLM infer the specific output format expected for a query. Whereas for other tasks like SST-2, SST-5, MNLI, MedQA, TREC, FinQA and GSM-8k (final answer-correctness) that do not have complex output formats, model performance is not sensitive to choice of ICE (or can be recovered with better instructions) and we see a high NICE score.
@@ 7 Conclusion
Our paper challenges the existing consensus that optimizing in-context examples (ICE) is universally beneficial. Through extensive experimentation on LLMs, we demonstrate that ICE optimization yields diminishing returns in the presence of task-specific instructions. We introduce a novel and efficient metric, NICE, to gauge the utility of optimizing ICE for a task. NICE measures the invariability of task performance with respect to the choice of ICE. This helps in efficient utilization of data and compute resources. A NICE score near 1 for a task implies minimal impact of ICE choice on model performance, allowing the use of static ICE that avoids query-dependent retrievals at inference time and improves computational efficiency. Thus for certain tasks, we show that a given LLM API budget may be more judiciously used to optimize the prompt instruction than to finetune available retriever models.
While we focus on using NICE to evaluate the effectiveness of ICE optimization for a task, the same can also be used to compare different instructions by optimizing a weighted sum of NICE scores (see more details in Appendix A.2.1). For simplicity, we use uniform weights in this work.
8 Limitations
We work with current state-of-the-art LLMs (GPT-4-Turbo, GPT-3.5-Turbo, Mixtral-8x7B-Instruct & Llama-2-70B-Chat) and the observations could differ with future models. Due to cost considerations, we do not compare more advanced learning-based methods for instruction and ICE optimization. While we do look at several tasks, we do not validate our hypotheses on completely new or open-ended generation tasks. Lastly, since training data for most models is not publicly available, data leakage from the training data corpus could possibly affects our findings.
@@ A.1.1 Classification Tasks
In 3.1, we discuss the structured instruction sets we use. In this section, we give some concrete examples. We also note that all of these instructions follow standard templates and none of them required human creativity. The total amount of time it took to create all of them for $9$ datasets was under $1$ hour. We show the instruction sets for SST-2 and MTOP, and refer to our codebase for all other datasets.
@@ A.1.2 Generation Tasks
3. The standard format of the output is:- [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]...[SL: <slot-label-n> <slot-label-value-n> ] ] if there are n slot labels tagged in an utterance.
4. In some cases, the slots can be nested with intents within them, for those cases, use this format:- [IN: <user-intent> [SL: <slot-label> [IN: <user-intent> [SL: <slot-label-1> <slot-label-value-1> ] [SL: <slot-label-2> <slot-label-value-2> ]...[SL: <slot-label-n> <slot-label-value-n>]]]]
@@
The ‘‘No Instruction’’ case for both classification and generation tasks has an empty instruction appended by the examples. For instruction templates for all our classification and generation tasks, we refer the interested reader to our codebase.